please explain how to solve: if p is 64% of q, then q is what percent of p?
Answer by nyc_function(2741)

If p = 0.64q, then q = xp?

p = 0.64q

p/0.64 = q

===========================

p/0.64 = x(0.64q)

p/0.64 divided by 0.64q = x

0.64q/0.64 times 1/0.64q = x

1/0.64 = x

1.5625 = x

156.25% = x